14. Suppose that on January 1, 2000, you purchased 1000 shares of MK-Soft stocks at $25 per share, and sold them on January 1, 2010 for $120.

A) What is your annual rate of return for the 10-year period?

B) What is your total rate of return for the 10-year period?

C) What if the stock has paid a $1.00 dividend annually for the 10-year period?
Answer #1

(a) Number of shares = 1000

Total Purchase Price = P = 1000*25 = $25000

Total Selling Price = F = 1000*120 = $120000

Number of Years = n

Annual Rate of Return = (F/P)1/n - 1 = (120000/25000)1/5 - 1 = 0.3685 or 36.85%

(b) Total Return for 10 Year period = F/P - 1 = 120000/25000 - 1 = 3.8 or 380%

(c) Dividend Paid in 10 Years = D = 1*1000*10 = $10000

Annual Rate of Return = ((F+D)/P)1/n - 1 = ((120000+10000)/25000)1/5 - 1 = 0.3905 or 39.05%

Total Return for 10 Year period = (F+D)/P - 1 = (120000+10000)/25000 - 1 = 4.2 or 420%